OpenStudy (anonymous):

Which of (3,5), (4,6), (5,7) and (6,8) are solutions to y=x+2 (3,5) and (4,6) (4,6) and (5,7) none all

OpenStudy (cgreenwade2000):

It's as easy as plugging in the x and y values and seeing if both asides are equal.

OpenStudy (anonymous):

to check if (3,5) is a solution for example you have to plug it into y = x+2 5 = 3 +2 5 =5 so (3,5) is a solution can you try and check if (4,6) is also a solution ?